OpenStudy (anonymous):

Find S22 for -6 + (-4) + (-2) + 0 … A. 330 B. 332 C. 334 D. 336

OpenStudy (anonymous):

c.d= d=-4-(-6)=-4+6=2 sn=n/1{2a+(n-1)d} n=22 a=-6 plug in the values and get the answer

OpenStudy (anonymous):

i keep getting 66

OpenStudy (anonymous):

\[S22=\frac{ 22 }{ 2 }\left\{ 2\left( -6 \right)+\left( 22-1 \right)2 \right\}=11\left( -12+42 \right)=330\]